Find the volume of a sphere that has a diameter of 9 meters.
A) 243 m3 B) 486 m3 C) 121.5 m3 D) 972 m3

Answer by feliz1965(151)

We need to use the volume of a sphere formula.
The formula is:
V = 4/3 times pi times r^3
You were given the dimater to be 9 meters.
The radius = 1/2 of the given diameter.
So, our radius is 4.5 meters.
We now plug and chug.
V = 4/3 times 3.14 times (4.5)^3
V = 1.3 times 3.14 times 91.125
Can you finish now?
By the way, the choices given do not match the right answer for the volume of this sphere. Are you sure there is no typo here?